OpenStudy (anonymous):

solve: ln(x^2) - ln 7=0

OpenStudy (slaaibak):

easiest method would be to say: \[\ln({x^2 \over 7}) = 0\] and we know the ln 1 = 0 so x^2 / 7 = 1 \[x = \pm \sqrt7\]

OpenStudy (anonymous):

thank you thank you thank you!

OpenStudy (slaaibak):

it's my pleasure :)